OpenStudy (anonymous):

Factor the polynomial: z4 - 10z2 + 9

OpenStudy (anonymous):

assume x=z^2 then your expression becomes x^2 -10x +9 (x-9)(x-1) (z^2-9)(z^2-1) a^2-b^2=(a-b)(a+b) (z-3)(z+3)(z-1)(z+1)